OpenStudy (anonymous):

in parallelogram DEFG, DH = x+1 HF= 3y , GH=3x-4 and HE+5y+1 find the value of x and y @sprinkles_reven

OpenStudy (anonymous):

here

OpenStudy (anonymous):

ok so 3x-4=x+1

OpenStudy (anonymous):

2x=6

OpenStudy (anonymous):

and 5y+1=3y

OpenStudy (anonymous):

where the 6 come frm

OpenStudy (anonymous):

oops i meant 5

OpenStudy (anonymous):

2x=5

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

so x = 2.5

OpenStudy (anonymous):

and 5y+1=3y

OpenStudy (anonymous):

2y=-1

OpenStudy (anonymous):

-2? or 2

OpenStudy (anonymous):

it depends on which side you subtract

OpenStudy (anonymous):

but i always like to have a positive variable

OpenStudy (anonymous):

makes things easier

OpenStudy (anonymous):

so y=-0.5

OpenStudy (anonymous):

okay okay ii get what your saying

OpenStudy (anonymous):

you good or do you need more help??

OpenStudy (anonymous):

remember to best response

OpenStudy (anonymous):

that would be great if you would fan me @south834